# Diffraction by a Right-Angle Penetrable Wedge: Closed-Form Solution for $\nu=\sqrt{2}$

### v2

@@ 7 Half-period shift identities at $e_{2}=-1$
$W0_{\ell}:=\wp(u_{0}-u_{\ell})=\frac{i}{t_{\ell}^{2}},\qquad W1_{\ell}:=\wp^{\prime}(u_{0}-u_{\ell})=\frac{(1-i)Y_{\ell}}{t_{\ell}^{3}},\qquad W2_{\ell}:=\wp^{\prime\prime}(u_{0}-u_{\ell})=-2-\frac{6}{t_{\ell}^{4}}.$ (42)
@@ 7.1 Specialization of the addition theorem
Using the differential equation for $(g_{2},g_{3})=(4,0)$ (see, e.g., [14, § 23.6]),
@@
$\frac{1}{4}\left(\frac{\wp^{\prime}(u)}{\wp(u)+1}\right)^{2}=\frac{\wp(u)\left(\wp(u)^{2}-1\right)}{(\wp(u)+1)^{2}}=\frac{\wp(u)(\wp(u)-1)}{\wp(u)+1}.$ (46)
@@
$\boxed{\ \wp^{\prime}(u+u_{0})=-\frac{2\wp^{\prime}(u)}{(\wp(u)+1)^{2}}\ }.$ (48)
Differentiating (48) yields
$\boxed{\ \wp^{\prime\prime}(u+u_{0})=-2\left(\frac{\wp^{\prime\prime}(u)}{(\wp(u)+1)^{2}}-\frac{2\wp^{\prime}(u)^{2}}{(\wp(u)+1)^{3}}\right)\ }.$ (49)
7.2 Specialization to $u_{0}-u_{\ell}$ and algebraic elimination via the uniformization
@@
$\wp(-u_{\ell})=\wp(u_{\ell}),\qquad\wp^{\prime}(-u_{\ell})=-\wp^{\prime}(u_{\ell}),\qquad\wp^{\prime\prime}(-u_{\ell})=\wp^{\prime\prime}(u_{\ell}),$ (50)
@@
$\wp(u_{0}-u_{\ell})=-1+\frac{2}{\wp(u_{\ell})+1}=\frac{1-\wp(u_{\ell})}{1+\wp(u_{\ell})},\qquad\wp^{\prime}(u_{0}-u_{\ell})=\frac{2\wp^{\prime}(u_{\ell})}{(\wp(u_{\ell})+1)^{2}}.$ (51)
Now substitute the uniformization $\wp(u_{\ell})=x_{W}(t_{\ell})$ and $\wp^{\prime}(u_{\ell})=2y_{W}(t_{\ell},Y_{\ell})$. A direct simplification gives
$\wp(u_{0}-u_{\ell})=\frac{i}{t_{\ell}^{2}},\qquad\wp^{\prime}(u_{0}-u_{\ell})=\frac{(1-i)Y_{\ell}}{t_{\ell}^{3}}.$ (52)
Finally, using $\wp^{\prime\prime}(u)=6\wp(u)^{2}-g_{2}/2=6\wp(u)^{2}-2$ for $g_{2}=4$,
$\wp^{\prime\prime}(u_{0}-u_{\ell})=6\left(\frac{i}{t_{\ell}^{2}}\right)^{2}-2=-2-\frac{6}{t_{\ell}^{4}},$ (53)
which proves (42).
@@ 8.1 Definitions
Let $\zeta_{W}(u)$ denote the Weierstrass zeta function, characterized by $\zeta_{W}^{\prime}(u)=-\wp(u)$ and $\zeta_{W}(u)\sim u^{-1}$ as $u\to 0$; see [14, § 23.2,§ 23.9] or [16, Ch. 20]. Define
$A(u)=\sum_{\ell\in I_{\mathrm{scat}}}\alpha_{\ell}\Bigl[\zeta_{W}(u-u_{\ell})-\zeta_{W}(u_{0}-u_{\ell})\Bigr],\qquad B(u)=\sum_{\ell\in I_{\mathrm{scat}}}\beta_{\ell}\Bigl[\zeta_{W}(u-u_{\ell})-\zeta_{W}(u_{0}-u_{\ell})\Bigr].$ (54)
@@
$p_{1}$ $=e^{-i\pi/4}\sum_{\ell\in I_{\mathrm{scat}}}\alpha_{\ell}\,W0_{\ell},\qquad p_{2}=e^{-i\pi/2}\frac{1}{2}\sum_{\ell\in I_{\mathrm{scat}}}\alpha_{\ell}\,W1_{\ell},\qquad p_{3}=e^{-3i\pi/4}\frac{1}{6}\sum_{\ell\in I_{\mathrm{scat}}}\alpha_{\ell}\,W2_{\ell},$ (56)
$q_{1}$ $=e^{-i\pi/4}\sum_{\ell\in I_{\mathrm{scat}}}\beta_{\ell}\,W0_{\ell},\qquad q_{2}=e^{-i\pi/2}\frac{1}{2}\sum_{\ell\in I_{\mathrm{scat}}}\beta_{\ell}\,W1_{\ell},\qquad q_{3}=e^{-3i\pi/4}\frac{1}{6}\sum_{\ell\in I_{\mathrm{scat}}}\beta_{\ell}\,W2_{\ell}.$ (57)
8.2 Local relation $u-u_{0}=e^{-i\pi/4}t+O(t^{5})$ (coefficient matching)
Let $\delta:=u-u_{0}$. From the uniformization $\wp(u)=x_{W}(t)$, expand $x_{W}(t)$ as $t\to 0$:
$x_{W}(t)=\frac{it^{2}+1}{it^{2}-1}=-1-2it^{2}+2t^{4}+O(t^{6}).$ (58)
@@ 8.2 Local relation $u-u_{0}=e^{-i\pi/4}t+O(t^{5})$ (coefficient matching)
Equating with $x_{W}(t)$ yields
$2\delta^{2}-2\delta^{4}=-2it^{2}+2t^{4}+O(t^{6}).$ (61)
Write $\delta^{2}=-it^{2}+at^{4}+O(t^{6})$ so that $\delta^{4}=(-it^{2})^{2}+O(t^{6})=-t^{4}+O(t^{6})$. Matching coefficients forces $a=0$, hence
$\delta=e^{-i\pi/4}t+O(t^{5}),$ (62)
@@ 8.3 Jet cancellation
$A(u)=-\delta\sum_{\ell}\alpha_{\ell}W0_{\ell}-\frac{\delta^{2}}{2}\sum_{\ell}\alpha_{\ell}W1_{\ell}-\frac{\delta^{3}}{6}\sum_{\ell}\alpha_{\ell}W2_{\ell}+O(\delta^{4}).$ (64)
Using (62) gives $\delta^{n}=e^{-in\pi/4}t^{n}+O(t^{n+4})$ for $n=1,2,3$, hence
$A(u(t))$ $=-e^{-i\pi/4}t\sum_{\ell}\alpha_{\ell}W0_{\ell}-e^{-i\pi/2}\frac{t^{2}}{2}\sum_{\ell}\alpha_{\ell}W1_{\ell}-e^{-3i\pi/4}\frac{t^{3}}{6}\sum_{\ell}\alpha_{\ell}W2_{\ell}+O(t^{4}).$ (65)
By definition of $(p_{1},p_{2},p_{3})$, the polynomial $p(t)$ is the negative of the displayed cubic truncation, so
$A(u(t))+p(t)=O(t^{4}),\qquad t\to 0\ \text{on}\ \Omega_{+}^{\mathrm{phys}}.$ (66)
@@ 9 $\tau^{2}$ pairing compression
$\ell=(m,\sigma,\varepsilon_{w})\longmapsto\ell^{\prime}=(m+2\bmod 4,\sigma,\varepsilon_{w}).$ (68)
Under this pairing, the pole transport implies
@@
Applying the derivative convention (30) under $t\mapsto-t$ yields $w^{\prime}_{m+2}(u_{\ell^{\prime}})=-w^{\prime}_{m}(u_{\ell})$ and hence
@@
Using the half-period identities (42), we have $W0_{\ell^{\prime}}=W0_{\ell}$, $W2_{\ell^{\prime}}=W2_{\ell}$, but $W1_{\ell^{\prime}}=-W1_{\ell}$. Moreover, from the coefficient tables in Proposition 6.16.2, the phase flip and $r_{I}$ flip cancel, so $\alpha_{\ell^{\prime}}=\alpha_{\ell}$ and $\beta_{\ell^{\prime}}=\beta_{\ell}$.
@@
  • Pair contributions double in $p_{1},p_{3},q_{1},q_{3}$ (built from $W0$ and $W2$).
  • Pair contributions cancel in $p_{2},q_{2}$ (built from $W1$).
@@
$p_{2}=p_{2}^{(\ell^{\prime}_{\mathrm{inc}})},\qquad q_{2}=q_{2}^{(\ell^{\prime}_{\mathrm{inc}})}.$ (72)
10 Singular channel, explicit $d_{\ell}$ table, and canonical no-double-counting theorem
@@ 10.1 Singular channel and residue definition
$\beta_{\mathrm{ch}}(t,Y):=\frac{iY}{\sqrt{2}}(t^{2}-1),$ (73)
@@
10.2 Explicit $d_{\ell}$ parity $\times j$ table
Substitute Proposition 6.16.2 into (75) and simplify using only $Y^{2}=2(t^{4}+1)$.
@@ 10.2 Explicit $d_{\ell}$ parity $\times j$ table
Write $(t,Y)=(t_{\ell},Y_{\ell})$ and $r_{I}=r_{I}(\ell)$.
$\text{If $m$ is even:}\quad d_{\ell}$ $=\begin{cases}\displaystyle\frac{\chi_{m}r_{I}}{4t^{2}}\bigl(t^{6}-t^{4}+t^{2}-2\bigr),&j=1,\\[5.16663pt] \displaystyle\frac{\chi_{m}r_{I}}{4t^{4}}\bigl(t^{6}-2t^{4}+2t^{2}-2\bigr),&j=2,\\[5.16663pt] \displaystyle\frac{\chi_{m}r_{I}}{4}\,\frac{iY}{\sqrt{2}}(t^{2}-1),&j=3,\\[5.16663pt] 0,&j=4;\end{cases}$
$\text{If $m$ is odd:}\quad d_{\ell}$ $=\begin{cases}\displaystyle\frac{\psi_{m}r_{I}}{4},&j=1,\\[4.30554pt] \displaystyle\frac{\psi_{m}r_{I}}{4}\,t^{4},&j=2,\\[4.30554pt] \displaystyle\frac{\kappa_{m}r_{I}}{4}\,\frac{Y}{\sqrt{2}}\,\frac{2-t^{2}}{t^{2}},&j=3,\\[5.16663pt] \displaystyle\frac{\kappa_{m}r_{I}}{4}\,\frac{Y}{\sqrt{2}}\,\frac{t^{6}-2}{t^{4}},&j=4.\end{cases}$
@@
In the cases (meven,j=1)(m\ $\text${even},j=1), (meven,j=2)(m\ $\text${even},j=2), and (modd,j=2)(m\ $\text${odd},j=2), the $\beta_{\ell}$ carries a factor $Y$, so $\beta_{\mathrm{ch}}(t,Y)\beta_{\ell}$ carries $Y^{2}$. Replacing $Y^{2}$ by $2(t^{4}+1)$ via $Y^{2}=2(t^{4}+1)$ removes $Y$, making these $d_{\ell}$ purely $t$-rational, as visible in Proposition 10.1.
@@ 10.3 Remainder $R(u)$ and pole cancellation
$R(u):=P_{1,3}(u)-\sum_{\ell\in I_{\mathrm{scat}}}d_{\ell}\Bigl[\zeta_{W}(u-u_{\ell})-\zeta_{W}(u_{0}-u_{\ell})\Bigr].$ (76)
@@
(i) $R(u)$ has no pole at any $u=u_{\ell}$, $\ell\in I_{\mathrm{scat}}$.
(ii) $R(u)$ is analytic at $u=u_{0}$ (equivalently at $t=0$ on $\Omega_{+}^{\mathrm{phys}}$).
@@
(i) By definition (75), the principal part of $P_{1,3}$ at $u_{\ell}$ is $d_{\ell}/(u-u_{\ell})$. The zeta difference $\zeta_{W}(u-u_{\ell})-\zeta_{W}(u_{0}-u_{\ell})$ has principal part $1/(u-u_{\ell})$. Subtracting $d_{\ell}[\cdots]$ cancels the principal part at each $u_{\ell}$.
(ii) By jet cancellation, $A(u(t))+p(t)=O(t^{4})$ and $B(u(t))+q(t)=O(t^{4})$ as $t\to 0$, so each fraction in $P_{1,3}$ is analytic at $t=0$. Each zeta difference vanishes at $u=u_{0}$, hence is analytic there. Therefore $R$ is analytic at $u_{0}$. ∎

### v3

@@ 7 Half-period shift identities at e2 = -1
$W_{0\ell}$ $=\wp(u_{0}-u_{\ell})=-\frac{\sqrt{2}\,t_{\ell}^{2}}{Y_{\ell}+\sqrt{2}},$ (39)
$W_{1\ell}$ $=\wp^{\prime}(u_{0}-u_{\ell})=-\frac{4t_{\ell}}{Y_{\ell}+\sqrt{2}},$
$W_{2\ell}$ $=\wp^{\prime\prime}(u_{0}-u_{\ell})=\frac{12t_{\ell}^{4}}{(Y_{\ell}+\sqrt{2})^{2}}-2.$
@@ 7.1 Specialization of the addition theorem
Using the differential equation for (20) (see, e.g., [14, §23.6]),
@@
$\frac{1}{4}\left(\frac{\wp^{\prime}(u)}{\wp(u)+1}\right)^{2}=\frac{\wp(u)\bigl(\wp(u)^{2}-1\bigr)}{(\wp(u)+1)^{2}}=\frac{\wp(u)(\wp(u)-1)}{\wp(u)+1}.$
@@
$\wp^{\prime}(u+u_{0})=-\frac{2\wp^{\prime}(u)}{(\wp(u)+1)^{2}}.$ (44)
Differentiating once more yields
$\wp^{\prime\prime}(u+u_{0})=-2\left(\frac{\wp^{\prime\prime}(u)}{(\wp(u)+1)^{2}}-\frac{2(\wp^{\prime}(u))^{2}}{(\wp(u)+1)^{3}}\right).$ (45)
7.2 Specialization to u0 - u_l and algebraic elimination
@@
$\wp(-u_{\ell})=\wp(u_{\ell}),\qquad\wp^{\prime}(-u_{\ell})=-\wp^{\prime}(u_{\ell}),\qquad\wp^{\prime\prime}(-u_{\ell})=\wp^{\prime\prime}(u_{\ell}),$
@@
$\wp(u_{0}-u_{\ell})=\frac{1-\wp(u_{\ell})}{1+\wp(u_{\ell})},\qquad\wp^{\prime}(u_{0}-u_{\ell})=\frac{2\wp^{\prime}(u_{\ell})}{(\wp(u_{\ell})+1)^{2}}.$ (46)
Now substitute the uniformization $\wp(u_{\ell})=x_{W}(t_{\ell},Y_{\ell})$ and $\wp^{\prime}(u_{\ell})=2y_{W}(t_{\ell},Y_{\ell})$ from (21). A direct simplification gives the first two identities in (39). Finally, using
$\wp^{\prime\prime}(u)=6\wp(u)^{2}-\frac{g_{2}}{2}=6\wp(u)^{2}-2\qquad(g_{2}=4),$
yields the third identity in (39). ∎
@@ 8.1 Definitions
Let $\zeta_{W}(u)$ denote the Weierstrass zeta function, characterized by $\zeta_{W}^{\prime}(u)=-\wp(u)$ and $\zeta_{W}(u)\sim u^{-1}$ as $u\to 0$. Define
$A(u)=\sum_{\ell\in I_{\mathrm{scat}}}\alpha_{\ell}\big[\zeta_{W}(u-u_{\ell})-\zeta_{W}(u_{0}-u_{\ell})\big],\qquad B(u)=\sum_{\ell\in I_{\mathrm{scat}}}\beta_{\ell}\big[\zeta_{W}(u-u_{\ell})-\zeta_{W}(u_{0}-u_{\ell})\big].$ (47)
@@
$p_{1}=\frac{1}{\sqrt{2}}\sum_{\ell\in I_{\mathrm{scat}}}\alpha_{\ell}W_{0\ell},\qquad p_{2}=\frac{1}{4}\sum_{\ell\in I_{\mathrm{scat}}}\alpha_{\ell}W_{1\ell},\qquad p_{3}=\frac{1}{12\sqrt{2}}\sum_{\ell\in I_{\mathrm{scat}}}\alpha_{\ell}W_{2\ell},$ (49)
$q_{1}=\frac{1}{\sqrt{2}}\sum_{\ell\in I_{\mathrm{scat}}}\beta_{\ell}W_{0\ell},\qquad q_{2}=\frac{1}{4}\sum_{\ell\in I_{\mathrm{scat}}}\beta_{\ell}W_{1\ell},\qquad q_{3}=\frac{1}{12\sqrt{2}}\sum_{\ell\in I_{\mathrm{scat}}}\beta_{\ell}W_{2\ell}.$ (50)
8.2 Local relation between u and t near the basepoint
Let $\delta:=u-u_{0}$. From the uniformization $\wp(u)=x_{W}(t,Y)$ and the physical branch $Y\sim-\sqrt{2}$ as $t\to 0$, expand $x_{W}(t,Y)$ as $t\to 0$ on $\Omega_{\mathrm{phys}}^{+}$:
$x_{W}(t,Y)=-1+t^{2}-\frac{1}{2}t^{4}+O(t^{6}).$ (51)
@@ 8.2 Local relation between u and t near the basepoint
Equating (51) and (53) yields
$2\delta^{2}-2\delta^{4}=t^{2}-\frac{1}{2}t^{4}+O(t^{6}).$
Writing $\delta^{2}=\frac{1}{2}t^{2}+at^{4}+O(t^{6})$ gives $\delta^{4}=\frac{1}{4}t^{4}+O(t^{6})$ and forces $a=0$, hence
$\delta=\frac{t}{\sqrt{2}}+O(t^{5}),$ (54)
@@ 8.3 Jet cancellation
$A(u)=-\delta\sum_{\ell}\alpha_{\ell}W_{0\ell}-\frac{\delta^{2}}{2}\sum_{\ell}\alpha_{\ell}W_{1\ell}-\frac{\delta^{3}}{6}\sum_{\ell}\alpha_{\ell}W_{2\ell}+O(\delta^{4}).$
Using (54) gives $\delta^{n}=(t/\sqrt{2})^{n}+O(t^{n+4})$ for $n=1,2,3$, hence
$A(u(t))=-\frac{t}{\sqrt{2}}\sum_{\ell}\alpha_{\ell}W_{0\ell}-\frac{t^{2}}{4}\sum_{\ell}\alpha_{\ell}W_{1\ell}-\frac{t^{3}}{12\sqrt{2}}\sum_{\ell}\alpha_{\ell}W_{2\ell}+O(t^{4}).$
By definition of $(p_{1},p_{2},p_{3})$ in (49), the polynomial $p(t)$ is the negative of the displayed cubic truncation, so
$A(u(t))+p(t)=O(t^{4}),\qquad t\to 0\ \text{on }\Omega_{\mathrm{phys}}^{+}.$ (56)
@@ 9 Tau-squared pairing compression
$\ell=(m,\sigma,\varepsilon_{w})\longmapsto\ell^{\prime}=(m+2\!\!\!\!\pmod{4},\ \sigma,\varepsilon_{w}).$ (58)
Under this pairing, the pole transport (31) implies
@@
Applying (33) under $t\mapsto-t$ yields $w^{\prime}_{m+2}(u_{\ell^{\prime}})=-w^{\prime}_{m}(u_{\ell})$ and hence
@@
Using (39), we have $W_{0\ell^{\prime}}=W_{0\ell}$, $W_{2\ell^{\prime}}=W_{2\ell}$, but $W_{1\ell^{\prime}}=-W_{1\ell}$. Moreover, from Propositions 6.16.2, the phase flip and $r_{I}$ flip cancel, so $\alpha_{\ell^{\prime}}=\alpha_{\ell}$ and $\beta_{\ell^{\prime}}=\beta_{\ell}$.
@@
  • Pair contributions double in $p_{1},p_{3},q_{1},q_{3}$ (built from $W_{0}$ and $W_{2}$).
  • Pair contributions cancel in $p_{2},q_{2}$ (built from $W_{1}$).
@@
$p_{2}=p^{(\ell^{\prime}_{\mathrm{inc}})}_{2},\qquad q_{2}=q^{(\ell^{\prime}_{\mathrm{inc}})}_{2}$. (62)
10 Singular channel, explicit d_l table, and canonical no-double-counting theorem
@@ 10.1 Singular channel and residue definition
$\beta_{\mathrm{ch}}(t,Y):=\frac{\mathrm{i}Y}{\sqrt{2}}\,(t^{2}-1),$ (63)
@@
10.2 Explicit d_l parity-by-j table
Substitute Propositions 6.16.2 into (65) and simplify using only $Y^{2}=2(t^{4}+1)$.
@@ 10.2 Explicit d_l parity-by-j table
Write $(t,Y)=(t_{\ell},Y_{\ell})$ and $r_{I}=r_{I}(\ell)$. If $m$ is even:
$d_{\ell}=\begin{cases}\displaystyle\chi_{m}\frac{r_{I}}{4t^{2}}\,(t^{6}-t^{4}+t^{2}-2),&j=1,\\[6.0pt] \displaystyle\chi_{m}\frac{r_{I}}{4t^{4}}\,(t^{6}-2t^{4}+2t^{2}-2),&j=2,\\[6.0pt] \displaystyle\chi_{m}\frac{r_{I}}{4}\,\frac{\mathrm{i}Y}{\sqrt{2}}(t^{2}-1),&j=3,\\[6.0pt] 0,&j=4;\end{cases}$
If $m$ is odd:
$d_{\ell}=\begin{cases}\displaystyle\psi_{m}\frac{r_{I}}{4},&j=1,\\[6.0pt] \displaystyle\psi_{m}\frac{r_{I}}{4}\,t^{4},&j=2,\\[6.0pt] \displaystyle\kappa_{m}\frac{r_{I}}{4}\,\frac{Y}{\sqrt{2}}\,\frac{2-t^{2}}{t^{2}},&j=3,\\[10.0pt] \displaystyle\kappa_{m}\frac{r_{I}}{4}\,\frac{Y}{\sqrt{2}}\,\frac{t^{6}-2}{t^{4}},&j=4.\end{cases}$
@@
In the cases (even $m$, $j=1$), (even $m$, $j=2$), and (odd $m$, $j=2$), the coefficient $\beta_{\ell}$ carries a factor $Y$, so $\beta_{\mathrm{ch}}(t,Y)\beta_{\ell}$ carries $Y^{2}$. Replacing $Y^{2}$ by $2(t^{4}+1)$ via $Y^{2}=2(t^{4}+1)$ removes $Y$, making these $d_{\ell}$ purely $t$-rational, as visible in Proposition 10.1.
@@ 10.3 Remainder R(u) and pole cancellation
$R(u):=P_{1,3}(u)-\sum_{\ell\in I_{\mathrm{scat}}}d_{\ell}\big[\zeta_{W}(u-u_{\ell})-\zeta_{W}(u_{0}-u_{\ell})\big].$ (66)
@@
  1. (i)
    $R(u)$ has no pole at any $u=u_{\ell}$, $\ell\in I_{\mathrm{scat}}$.
  2. (ii)
    $R(u)$ is analytic at $u=u_{0}$ (equivalently at $t=0$ on $\Omega_{\mathrm{phys}}^{+}$).
@@
(i) By definition (65), the principal part of $P_{1,3}$ at $u_{\ell}$ is $d_{\ell}/(u-u_{\ell})$. The zeta difference $\zeta_{W}(u-u_{\ell})-\zeta_{W}(u_{0}-u_{\ell})$ has principal part $1/(u-u_{\ell})$. Subtracting $d_{\ell}[\cdot]$ cancels the principal part at each $u_{\ell}$.
(ii) By jet cancellation (56)–(57), $A(u(t))+p(t)=O(t^{4})$ and $B(u(t))+q(t)=O(t^{4})$ as $t\to 0$, so each fraction in (64) is analytic at $t=0$. Each zeta difference vanishes at $u=u_{0}$, hence is analytic there. Therefore $R$ is analytic at $u_{0}$. ∎